OpenStudy (anonymous):

Find an equation of the plane consisting of all points that are equidistant from (1, 3, -4) and (2, -3, 1), and having 1 as the coefficient of x

OpenStudy (anonymous):

last one i swear

OpenStudy (amistre64):

find the vector between the 2 points and the distance/2

OpenStudy (amistre64):

distance is subtract and pythag (1, 3, -4) -(2, -3, 1) ----------- -1 6 -5 ; sqrt(1+36+25) = sqrt(62) right?

OpenStudy (amistre64):

and the subtracted part is the vector for the normal we can use

OpenStudy (amistre64):

<1,-6, 5> = normal vector to find a midpoint ... add the points and divide by 2 lol

OpenStudy (anonymous):

how u retain all this information truly remarkable

OpenStudy (amistre64):

(1, 3, -4) (2, -3, 1) ----------- (3 , 0 ,-3)/2

OpenStudy (amistre64):

we use this point and the normal ... thatll get us all we need :)

OpenStudy (anonymous):

how do i set this equal to 0 then?

OpenStudy (amistre64):

1(x-(3/2))-6(y-0)+5(z-(-3/2))=0

OpenStudy (anonymous):

u happen to know how to determine a point of intersection given two lines?

OpenStudy (amistre64):



OpenStudy (amistre64):

i can recall some of it :) you have to parametric them and change the scalar to see if they cross

OpenStudy (amistre64):

the best view i got is that vectors/lines in space can move past each other and never touch; like a floor line and a ceiling line

OpenStudy (amistre64):

so to see where they are at any given time we change the scalar for the parametrics and equate them